OpenStudy (anonymous):

A pick-up truck is fitted with new tires which have a diameter of 42 inches. How fast will the pick-up truck be moving when the wheels are rotating at 280 revolutions per minute? Express the answer in miles per hour rounded to the nearest whole number.

OpenStudy (anonymous):

\[\frac{42\pi *280*60}{5280}=420 \text{ mph} \]

OpenStudy (anonymous):

42π∗280∗605280=420 mph